OpenStudy (anonymous):

Amaya height is 2 inches less than 2\5 the depth of the pool,d.if amaya is 52 inches tall how many deep is the pool? Part A: Write an equation that represents th problem solve the equation show your work Answer__________inches

OpenStudy (anonymous):

i need helpp ;'(

OpenStudy (dumbcow):

H=height of amaya D=depth of pool H = 2/5*D -2 H = 52 54 = 2/5*D -2 add 2 both sides 56 = 2/5*D multiply by 5/2 on both sides D = 28*5 D = 140

OpenStudy (anonymous):

thank you dumbcow:D

OpenStudy (anonymous):

Let p for the pool A) 52=2\5p-2 Solve: 2\5p =52+2 p= 54*(5/2), (time flip ) p= 135in The pool is 135 in

OpenStudy (dumbcow):

oops i had a typo follow what nancy said

OpenStudy (anonymous):

my is righ?

OpenStudy (anonymous):

thank you nancy also

OpenStudy (anonymous):

you welcome

OpenStudy (anonymous):

i got to go bye thank you guys